# SIC-POVMs and orders of real quadratic fields

### v1

@@ 12. Concluding Remarks
This paper considered fields that are geometric invariants of a SIC. It can be useful in practice to define SICs using data in smaller subfields that are not geometric invariants. One is the field generated by the entries of a single fiducial projector $\Pi$. That field can be smaller than the full projector field by nearly a factor of $d$ is some special cases, and this is extremely useful for computing exact SICs in some large dimensions that would otherwise be inaccessible to numerical investigation [6, 11]. Another is the field generated by the SIC overlaps $\operatorname{Tr}(\Pi D_{p})$ (or else the closely related SIC overlap phases $\sqrt{d+1}\operatorname{Tr}(\Pi D_{p})$), where $D_{p}$ are the displacement operators.
The overlap field (and the overlap phase field) appear to be ray class fields with ramification at just one of the infinite places $\infty_{1}$, provided one chooses a “strongly centered” fiducial projector; see [9, 20]. That phenomenon relates to the connection to Stark units, which live in a ray class field with ramification at the other infinite place $\infty_{2}$ and are (at least in some cases) Galois conjugate to squares of overlap phases.
Work in preparation by Appleby, Flammia, and Kopp [8] will make the classification scheme suggested by this paper explicit, giving a conjectural construction of each presumptive equivalence class of Weyl–Heisenberg covariant SICs in dimension $d$. Specifically, an explicit construction of a candidate for a map $\Omega$ satisfying the conditions of 8.5 (our strongest conjecture) will be given.
@@ 13. Acknowledgements
We thank Markus Grassl for finding numerical examples of degeneration of SIC-fields and sharing those examples with us. Professor Grassl’s insights prompted the first author to correct an originally incorrect statement and proof of 9.7.
We thank Marcus Appleby, Steven Flammia, and Gary McConnell for valuable conversations. We also acknowledge the contributions of Steve Donnelly, who independently connected SICs to orders in real quadratic number fields and shared his insights with Marcus Appleby, Steven Flammia, and Markus Grassl through informal discussions.
Appendix A Proof of 9.4
We will now prove 9.4, which we restate as A.1.
Lemma A.1.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta})$. Let $\Sigma$ be a subset of the two real embeddings of $\mathbb{Q}(\sqrt{\Delta})$. Fix an order $\mathcal{O}$ satisfying
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (A.1)
@@
$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (A.2)
@@
$\mathcal{O}_{K}=\left\{\frac{1}{2}(a+b\sqrt{D}):(a,b)\in\mathbb{Z}^{2},\,a\equiv b\,\left(\operatorname{mod}\,2\right),\,a\equiv 0\,\left(\operatorname{mod}\,2\right)\mbox{ if }D\not\equiv 1\,\left(\operatorname{mod}\,4\right)\right\}.$ (A.3)
$\operatorname{U}_{\mathfrak{m},\Sigma}(\mathcal{O}):=\left\{\alpha\in\mathcal{O}^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right),\,\sigma(\alpha)>0\mbox{ for }\sigma\in\Sigma\right\}.$ (A.4)
For A.1, the unit group of interest is $\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})$. The unit $\varepsilon_{d}$ is totally positive, and
$\varepsilon_{d}^{3}=\frac{1}{2}\left((d-1)(d^{2}-2d-2)+d(d-2)f\sqrt{D}\right)\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{\!\Delta}\right).$ (A.5)
Thus, we have the containment relations
$\langle\varepsilon_{d}^{3}\rangle\subseteq\operatorname{U}_{d\mathcal{O}_{\!\Delta},\{\infty_{1},\infty_{2}\}}\subseteq\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})\subseteq\operatorname{U}_{d\mathcal{O}_{K}}(\mathcal{O}_{K}).$ (A.6)
These relations show it suffices to prove (A.2) for the special case $\mathcal{O}=\mathcal{O}_{K}$ with $\Sigma=\emptyset$.
The proof studies the minimal unit $\eta>1\in\mathcal{O}_{K}$ such that
$\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.7)
@@
Claim 1. We have $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=1$. Consequently $\eta$ is totally positive, and $10<\eta^{\prime}<$ 1.
Proof. We know $\eta\eta^{\prime}=\pm 1$. By hypothesis, $\eta=\pm 1+d\omega$ for some $\omega\in\mathcal{O}_{K}$, and taking algebraic conjugates yields $\eta^{\prime}=\pm 1+d\omega^{\prime}$. Then $\eta\eta^{\prime}=1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, which forces $\eta\eta^{\prime}=1$. Total positivity and the bound for $\eta^{\prime}$ follow, proving the claim.
@@
Proof. We have $1<\eta\leq\varepsilon_{d}^{3}$, so for some $k\geq 1$, there holds $\eta^{k-1}<\varepsilon_{d}^{3}\leq\eta^{k}$. Then $\omega:=\eta^{k}\varepsilon_{d}^{-3}$ is a unit, $1\leq\omega<\eta$, and $\omega\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence by minimality of $\eta>1$, we must have $\omega=1$, proving the claim.
Claim 3. $\eta^{k}=\varepsilon_{d}^{3}$ for $k\geq 3$ cannot occur.
Proof. If $k=3$, then $\eta^{3}=\varepsilon_{d}^{3}$, which, since $\eta>1$ forces $\eta=\varepsilon_{d}$. We know $\eta\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and obtain a contradiction by showing that $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. To show the latter, we argue by contradiction and suppose $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. We know $\varepsilon_{d}\varepsilon_{d}^{\prime}=\eta\eta^{\prime}=1$, which then holds $\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and thus implies $\varepsilon_{d}^{\prime}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence, $\varepsilon_{d}+\varepsilon_{d}^{\prime}\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. However, $\varepsilon_{d}+\varepsilon_{d}^{\prime}=\frac{1}{2}(d-1+\sqrt{\Delta})+\frac{1}{2}(d-1+\sqrt{\Delta})=d-1\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, giving a contradiction for $d\geq 4$. Thus $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ holds, and $\eta\neq\varepsilon_{d}$.
If $k\geq 4$, then $\eta=(\varepsilon_{d})^{3/k}<\varepsilon_{d}$ (taking the positive $k$-th root.) Now Claim 1 gives (since $x+\frac{1}{x}$ is an increasing function for $x\geq 1$) that
$11<\eta+\eta^{\prime}=\operatorname{Tr}(\eta)<\varepsilon_{d}+\varepsilon_{d}^{\prime}=\operatorname{Tr}(\varepsilon_{d})=d-$ 1 . (A.8)
The trace is an integer, and
$\operatorname{Tr}(\eta)$ $\equiv\eta+\eta^{\prime}\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.9)
Since $d\geq 4$ and $\operatorname{Tr}(\eta)\in\mathbb{Z}$, the bounds on $\operatorname{Tr}(\eta)$ give
$\operatorname{Tr}(\eta)=\begin{cases}2&\mbox{if }\eta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right),\\ d-2&\mbox{if }\eta\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).\end{cases}$ (A.10)
@@
  • Case 2: Suppose $\operatorname{Tr}(\eta)=d-2$. We may assume $d\geq 5$, since the case $d=4$ is covered by Case 1. Now
    $d-3<\eta=\operatorname{Tr}(\eta)-\eta^{\prime}<\varepsilon_{d}<\operatorname{Tr}(\varepsilon_{d})=d-1.$ (A.11)
    We have for $d\geq 6$ that
    $(d-3)^{3}<\eta^{3}<\varepsilon_{d}^{3}<(d-1)^{3}<(d-3)^{5}<\eta^{5},$ (A.12)
    since $(d-1)^{3}<(d-3)^{5}$ for $d\geq 6$. Thus $\eta^{k}=\varepsilon_{d}^{3}$ cannot hold. For $d=5$, we have $\eta+\eta^{\prime}=3$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{3}-3x+1$, so $\eta=\frac{3+\sqrt{5}}{2}$. However, for $d=5$, $\Delta=12$, so $\varepsilon_{5}\in\mathbb{Q}(\sqrt{3})$ cannot be an integer power of $\eta$.
This completes the proof of Claim 3. It remains to rule out $k=2$, where $\varepsilon_{d}^{3}$ is the square of a unit $\tilde{\eta}$. We first classify all $d$ where such a unit exists.
@@
    $\eta=\varepsilon_{a+1}=\frac{a+\sqrt{a^{2}-4}}{2}$ (A.13)
@@
$\eta^{2}=\frac{1}{4}\left((a^{2}+b^{2}D)+2ab\sqrt{D}\right),$ (A.14)
while $\varepsilon_{d}=\frac{1}{2}(d-1+f\sqrt{D})$. To obtain a solution, we must have
$ab$ $=f;$ (A.15)
$a^{2}+b^{2}D$ $=2(d-1).$ (A.16)
We have $a\neq 0$ since $f\neq 0$, so the first equation gives $b^{2}=\frac{f^{2}}{a^{2}}$, and the second equation can be rewritten (since $\Delta=Df^{2}$) as
$a^{2}+\frac{\Delta}{a^{2}}=2(d-1).$ (A.17)
@@
  • Case 1: Suppose $d=a^{2}+1$ (so $a\geq 3$). Then $\Delta=a^{2}(a^{2}-4)$, $\sqrt{\Delta}=a\sqrt{a^{2}-4}$, and $a^{2}-4=Db^{2}$. Now
    $\eta$ $=\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4});$ (A.19)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}-2+a\sqrt{a^{2}-4}).$ (A.20)
@@
    $\eta$ $=\frac{1}{2}(a+\sqrt{a^{2}+4});$ (A.21)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}+2+a\sqrt{a^{2}+4}).$ (A.22)
@@
This proves Claim 4. To complete the proof, we must rule out the case $k=2$.
@@
We must show that all the $\tilde{\eta}$ described by Claim 4 have
$\tilde{\eta}\not\equiv\pm 1\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.23)
@@
$\tilde{\eta}=\frac{1}{2}\left(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}\right).$ (A.24)
Now $\Delta=(d+1)(d-3)=a^{2}(a^{2}-4)$; hence, $\Delta=f^{2}D$ has $D\mid(a^{2}-4)$ and $K=\mathbb{Q}(\sqrt{a^{2}-4})$. We assert for $d=a^{2}-1$ that
$\tilde{\eta}\equiv-a\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.25)
@@
    $\tilde{\eta}$ $=\frac{1}{2}(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}))$ (A.26)
    $\equiv\frac{1}{2}a(a^{2}-3)-\frac{1}{2}(a^{2}-1)$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.27)
    $\equiv\frac{a-1}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.28)
    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.29)

### v3

@@ 9. Concluding remarks
We state a conjectures refining 1.2 and 1.3. We also discuss other number fields attached to SICs that have been studied in the literature.
9.1. The Ideal-Class-to-SICs Conjecture
We formulate a final conjecture that refines 1.2 and 1.3 by insisting on compatibility with the Artin map. We formulate 9.1 because its truth would provide a unifying principle explaining the structure of all the other conjectures. In this paper, we present no mechanism to construct a candidate map $\Omega$ made in its statement.
To state this conjecture, we decompose
by means of 5.3.
Conjecture 9.1 (Ideal-Class-to-SIC Conjecture).
Fix a positive integer $d>3$, let $\Delta=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ where $\Delta_{0}$ is a fundamental discriminant of a quadratic field. Then there is a bijection $\Omega$,
$\Omega:\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\xrightarrow{\Omega}\operatorname{WHSIC}(d)/\operatorname{PEC}(d),$ (9.2)
Furthermore, $\Omega$ can be chosen to be compatible with the Artin map (defined in 6.2)
$\operatorname{Art}_{\mathcal{O}^{\prime}}:\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\to\operatorname{Gal}(\tilde{E}_{d,f^{\prime}}/\mathbb{Q}(\Delta_{0})),$ (9.3)
in the sense that
The required compatibility of $\Omega$ with the Artin map implies that there is exactly one Galois multiplet for each order $\mathcal{O}$ between $\mathcal{O}_{\!\Delta_{0}}$ and $\mathcal{O}_{\!\Delta}$.
9.1 would define a map $\mathcal{M}$ of the type given in 1.2 by means of the following commutative diagram,
9.1 is compatible with conjectures of the first author [33] and other authors [5, 10] connecting SICs to the Stark conjectures for real quadratic fields [43, 42], which posit the existence of special algebraic units constructed from partial zeta functions and carrying a Galois action compatible with the Artin map. Work in preparation by Appleby, Flammia, and Kopp [7] will make the classification scheme suggested by 9.1 explicit. The explicit classification also yields exact formulas for the stabilizers $\operatorname{stab}(S)$ and thus the size of the set $\operatorname{WHSIC}(d)$ discussed in 5.6.
9.2. Refined SIC fields
This paper considered fields that are geometric invariants of a SIC. It can be useful in practice to define SICs using data in smaller subfields that are not geometric invariants. One such subfield is the field generated by the entries of a single fiducial projector $\Pi$. That field can be smaller than the full projector field by nearly a factor of $d$ is some special cases, and this is extremely useful for computing exact SICs in some large dimensions that would otherwise be inaccessible to numerical investigation [5, 10]. Another is the field generated by the SIC overlaps $\operatorname{Tr}(\Pi D_{p})$ (or else the closely related SIC overlap phases $\sqrt{d+1}\operatorname{Tr}(\Pi D_{p})$), where $D_{p}$ are the displacement operators.
The overlap field (and the overlap phase field) both appear to be ray class fields specified by ramification at just one of the infinite places $\infty_{1}$, provided one chooses a “strongly centered” fiducial projector; see [8, 20]. That phenomenon relates to the connection to Stark units, which live in a ray class field with ramification at the other infinite place $\infty_{2}$ and are (at least in some cases) Galois conjugate to squares of overlap phases.
@@ 10. Acknowledgements
We thank Markus Grassl for finding numerical examples of degeneration of SIC fields and sharing those examples with us. We thank Marcus Appleby, Steven Flammia, and Gary McConnell for valuable conversations. We also acknowledge the contributions of Steve Donnelly, who independently connected SICs to orders in real quadratic number fields and shared his insights with Marcus Appleby, Steven Flammia, and Markus Grassl through informal discussions.
Appendix A Conjectures for triple product fields
We formulate conjectures about the triple product fields $F^{\rm trip}\!\left(S\right)$. Recall that $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$. We have limited empirical evidence for these conjectures.
Recall that $E_{d,f^{\prime}}=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}.$ Note that $E_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime}}$ if $d$ is odd, and $E_{d,f^{\prime}}\subseteq\tilde{E}_{d,f^{\prime}}$ if $d$ is even.
Conjecture A.1 (Extended Ray Class Fields of Orders Conjecture).
Fix an integer $d\geq 4$ and write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathbb{Q}(\sqrt{\Delta}))$. Then the bijection $\mathcal{M}$ of 1.3 can be chosen to satisfy the properties: If $S$ is a Weyl–Heisenberg SIC with $[S]\in\mathcal{M}(f^{\prime})$, then
$F^{\rm trip}\!\left(S\right)$ $=E_{d,f^{\prime}}$ and $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ $=\tilde{E}_{d,f^{\prime}}.$ (A.1)
A.1 implies part 1 of the following conjecture, which can be tested against data.
Conjecture A.2.
Let $d\geq 4$. Then for a Weyl–Heisenberg SIC $S$,
  • (1)
    The index
    $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]=\begin{cases}1&\mbox{ if }d\mbox{ is odd},\\ 2&\mbox{ if }d\mbox{ is even}.\end{cases}$ (A.2)
  • (2)
    $F^{\rm proj}\!\left(S\right)=F^{\rm trip}\!\left(S\right)(\xi_{d})$.
We have numerically verified part (1) of this conjecture, the index computation, for $4\leq d\leq 10$. Recall that 3.6(2) showed unconditionally that $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]$ divides $d^{2}$.
Appendix B Proof of 6.5
For completeness, this appendix presents results that are to appear elsewhere [36], reformulated using the variable $n=d+1$.
$\mathcal{O}_{K}=\left\{\frac{1}{2}(a+b\sqrt{D}):(a,b)\in\mathbb{Z}^{2},\,a\equiv b\,\left(\operatorname{mod}\,2\right),\,a\equiv 0\,\left(\operatorname{mod}\,2\right)\mbox{ if }D\not\equiv 1\,\left(\operatorname{mod}\,4\right)\right\}.$ (B.1)
$\operatorname{U}_{\mathfrak{m},\Sigma}(\mathcal{O}):=\left\{\alpha\in\mathcal{O}^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right),\,\sigma(\alpha)>0\mbox{ for }\sigma\in\Sigma\right\}.$ (B.2)
6.5 depends on the following lemma concerning unit groups of orders, which is special to the family $\Delta=(d+1)(d-3)$ of real quadratic orders, with $d\geq 4$.
Lemma B.1.
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (B.3)
@@ Appendix A Conjectures for triple product fields
$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (B.4)
For B.1, the unit group of interest is $\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})$. The unit $\varepsilon_{d}$ is totally positive, and
$\varepsilon_{d}^{3}=\frac{1}{2}\left((d-1)(d^{2}-2d-2)+d(d-2)f\sqrt{D}\right)\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{\!\Delta}\right).$ (B.5)
Thus, we have the containment relations
$\langle\varepsilon_{d}^{3}\rangle\subseteq\operatorname{U}_{d\mathcal{O}_{\!\Delta},\{\infty_{1},\infty_{2}\}}\subseteq\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})\subseteq\operatorname{U}_{d\mathcal{O}_{K}}(\mathcal{O}_{K}).$ (B.6)
These relations show it suffices to prove (B.4) for the special case $\mathcal{O}=\mathcal{O}_{K}$ with $\Sigma=\emptyset$.
The following stronger assertion in the special case, about the minimal unit $\eta\in\mathcal{O}_{K}$ such that $\eta>1$ and $\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, immediately implies Lemma B.1.
Lemma B.2.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta})$. Let $K=\mathbb{Q}(\sqrt{\Delta}).$ Then the smallest unit $\eta>1$ in $\mathcal{O}_{K}$ having
$\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (B.7)
is $\eta=\varepsilon_{d}^{3}$.
@@
Let $\Delta_{0}$ be the discriminant of the maximal order $\mathcal{O}_{K}$. Note that $\varepsilon_{d}>1$ is a unit of $\mathcal{O}_{K}$ of positive norm and is a root of $x^{2}-(d-1)x+1=0$. If $\varepsilon_{d}^{\prime}>0$ is its Galois conjugate, then $\varepsilon_{d}^{\prime}=\frac{1}{\varepsilon_{d}}$ and $\operatorname{Tr}(\varepsilon_{d})=\varepsilon_{d}+\varepsilon_{d}^{\prime}=d-1$, and $\varepsilon_{d}<d-1$.
@@
Claim 1. We have $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=1$. Then $\eta^{\prime}=\frac{1}{\eta}$ and $10<\eta^{\prime}<$ 1.
Proof. By hypothesis, $\eta=\pm 1+d\omega$ for some $\omega\in\mathcal{O}_{K}$, and taking Galois conjugates yields $\eta^{\prime}=\pm 1+d\omega^{\prime}$. It follows that $\eta\eta^{\prime}=1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. Now $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=\pm 1$ and $1\not\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ since $d\geq 4$, so we conclude $\eta\eta^{\prime}=1$. Thus $\eta^{\prime}>0$ and $10<\eta^{\prime}<$ 1, proving the claim.
@@
Proof. We have $1<\eta\leq\varepsilon_{d}^{3}$, hence there holds $\eta^{k-1}<\varepsilon_{d}^{3}\leq\eta^{k}$, for some $k\geq 1.$ Then $\omega:=\eta^{k}\varepsilon_{d}^{-3}$ is a unit, $1\leq\omega<\eta$, and $\omega\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. However, by minimality of $\eta>1$, we must have $\omega=1$, proving the claim.
Claim 3. $\eta^{k}=\varepsilon_{d}^{3}$ for some $k\geq 3$ cannot occur.
Proof. If $k=3$, then we have $\eta^{3}=\varepsilon_{d}^{3}$, and, since $\eta>1$, we have $\eta=\varepsilon_{d}$. By hypothesis $\eta\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, so we deduce $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. We rule out this possibility, arguing by contradiction. So suppose $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. We know $\varepsilon_{d}\varepsilon_{d}^{\prime}=\eta\eta^{\prime}=1$, which viewed $\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ implies $\varepsilon_{d}^{\prime}\equiv\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ We conclude that $\varepsilon_{d}+\varepsilon_{d}^{\prime}\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. On the other hand
$\varepsilon_{d}+\varepsilon_{d}^{\prime}=\frac{1}{2}(d-1+\sqrt{\Delta})+\frac{1}{2}(d-1+\sqrt{\Delta})=d-1\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (B.8)
Since ±$\pm 2\not\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ for any $d\geq 4$, we have a contradiction, whence it always holds that $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
If $k\geq 4$, then $\eta=(\varepsilon_{d})^{3/k}<\varepsilon_{d}$ (taking the positive $k$-th root.) We study the trace $\operatorname{Tr}(\eta)=\eta+\eta^{\prime}$ and note, by Claim 1, $\operatorname{Tr}(\eta)=\eta+\frac{1}{\eta}.$
Now, since $x+\frac{1}{x}$ is an increasing function for $x\geq 1$, and $1<\eta<e_{d}$, we have
$11<\eta+\eta^{\prime}=\operatorname{Tr}(\eta)<\varepsilon_{d}+\varepsilon_{d}^{\prime}=\operatorname{Tr}(\varepsilon_{d})=d-$ 1 . (B.9)
Now $\operatorname{Tr}(\eta)\in\mathbb{Z}$, and
$\operatorname{Tr}(\eta)$ $\equiv\eta+\eta^{\prime}\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (B.10)
Since $d\geq 4$, the bounds (B.9) on $\operatorname{Tr}(\eta)$ give
$\operatorname{Tr}(\eta)=\begin{cases}2&\mbox{if }\eta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right),\\ d-2&\mbox{if }\eta\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).\end{cases}$ (B.11)
@@
  • Case 2: Suppose $\operatorname{Tr}(\eta)=d-2$. In this case $\eta$ satisfies $x^{2}-(d-2)x+1=0$, hence $\eta=d-2-\eta^{\prime}>d-3$. For $d\geq 7$, noting that $(d-1)^{3}<(d-3)^{4}$ holds, and recalling $\varepsilon_{d}<d-1$, we have
    $(d-3)^{3}<\eta^{3}<\varepsilon_{d}^{3}<(d-1)^{3}<(d-3)^{4}<\eta^{4}\leq\eta^{k},$ (B.12)
    so there are no solutions $\eta^{k}=\varepsilon_{d}^{3}$ for $k\geq 4$ in this case.
    For the remaining values $d=4,5,6$, case $d=4$ is already ruled out in Case 1; $d=5$ has $\eta=\frac{3+\sqrt{5}}{2}$, while $\varepsilon_{5}=2+\sqrt{3}$; $d=6$ has $\eta=2+\sqrt{3}$ while $\varepsilon_{6}=\frac{5+\sqrt{21}}{2}$, so $\eta^{k}=\varepsilon_{d}^{3}$ has no solution since $\eta$ and $\varepsilon_{d}$ are in different number fields.
This completes the proof of Claim 3.
It remains to rule out $k=2$, where $\varepsilon_{d}^{3}$ is the square of a unit $\tilde{\eta}$. We first classify all $d$ where such a unit exists.
@@
    $\eta=\varepsilon_{a+1}=\frac{a+\sqrt{a^{2}-4}}{2}$ (B.13)
@@
$\eta^{2}=\frac{1}{4}\left((a^{2}+b^{2}\Delta_{0})+2ab\sqrt{\Delta_{0}}\right),$ (B.14)
while $\varepsilon_{d}=\frac{1}{2}(d-1+f\sqrt{\Delta_{0}})$. To obtain a solution, we must have
$ab$ $=f;$ (B.15)
$a^{2}+b^{2}\Delta_{0}$ $=2(d-1).$ (B.16)
We have $a\neq 0$ since $f\neq 0$, so the first equation gives $b^{2}=\frac{f^{2}}{a^{2}}$, and the second equation can be rewritten (since $\Delta=\Delta_{0}f^{2}$) as
$a^{2}+\frac{\Delta}{a^{2}}=2(d-1).$ (B.17)
@@
  • Case 1: Suppose $d=a^{2}+1$ (so $a\geq 3$). Then $\Delta=a^{2}(a^{2}-4)$, $\sqrt{\Delta}=a\sqrt{a^{2}-4}$, and $a^{2}-4=\Delta_{0}b^{2}$. Now
    $\eta$ $=\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4});$ (B.19)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}-2+a\sqrt{a^{2}-4}).$ (B.20)
@@
    $\eta$ $=\frac{1}{2}(a+\sqrt{a^{2}+4});$ (B.21)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}+2+a\sqrt{a^{2}+4}).$ (B.22)
@@
This proves Claim 4.
To complete the proof, we rule out the final case $k=2$.
@@
Proof. We must show that all the units $\tilde{\eta}$ described by Claim 4 have
$\tilde{\eta}\not\equiv\pm 1\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.23)
@@
$\tilde{\eta}=\frac{1}{2}\left(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}\right).$ (B.24)
Now $\Delta=(d+1)(d-3)=a^{2}(a^{2}-4)$; hence, $\Delta=f^{2}\Delta_{0}$ has $\Delta_{0}\mid(a^{2}-4)$ and $K=\mathbb{Q}(\sqrt{a^{2}-4})$. We assert for $d=a^{2}-1$ that
$\tilde{\eta}\equiv-a\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.25)
@@
    $\tilde{\eta}$ $=\frac{1}{2}(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}))$ (B.26)
    $\equiv\frac{1}{2}a(a^{2}-3)-\frac{1}{2}(a^{2}-1)$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (B.27)
    $\equiv\frac{a-1}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (B.28)
    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.29)